# Synergies Between Federated Learning and O-RAN: Towards an Elastic Architecture for Multiple Distributed Machine Learning Services

### v4

@@ III Case Study and Prototyping
For prototyping, we consider the dense urban O-RAN scenario in [8]. We assume $7$ O-RUs providing coverage across a hexagonal grid, with each O-RU having 1500 KHz of available bandwidth to allocate to FL clients. O-RAN environment, including signal forms, interference, channels, and data-rates are all modeled according to [9], where $32$ clients are distributed randomly across O-RUs, moving according to Lévy walk [10].
We consider two FL-services, each recruiting 16 clients, which utilize a convolutional neural network (CNN) to train on the MNIST (FL-service 1) and Fashion-MNIST (FL-service 2) datasets [11]. In FL-service 1, data points are distributed based on Dirichlet distribution [12] with parameter $\alpha{=}10$. For FL-service 2, we use Dirichlet distribution with $\alpha{=}0.01$, making clients’ datasets highly heterogeneous. Here, the parameter $\alpha$ controls the degree of data heterogeneity, with lower values creating greater heterogeneity. Each FL-service runs for 10 global training rounds. In each round, clients receive the global model and perform 30 local SGD iterations on their local data. Clients then send their trained local models back to the server, where global model aggregation is carried out using the FedAvg method [13]. We assume that FL-service 1 has a delay constraint of $30$ seconds to receive local models from its clients, while FL-service 2 has a stricter delay constraint of $20$ seconds. Clients who fail to completely transfer their local models to their assigned O-RUs within the designated time window are considered unsuccessful.
Figure 5: The impact of the importance of FL-service 2 in the objective function of near-RT FL controller (i.e., $a_{2}$) on the model prediction accuracy of FL-services. A higher value of $a_{2}$ leads to a shift of network resources (i.e., bandwidth) to the clients of FL-service 2, reducing its number of unsuccessful clients and improving its model prediction performance. Nevertheless, this comes with allocating less resources to FL-service 1, increasing its unsuccessful clients and degrading its model accuracy.
@@ III-A1 Non-RT system descriptor
To prototype this component, we first focus on arrows numbered 1 and 2 at the bottom of Fig. 2 and implement features for storing client descriptor (containing clients’ locations and dataset types) and FL-service descriptors (containing delay constraints for receiving local models). Considering the arrow numbered 3 in Fig. 2, we develop an eApp for mobility pattern prediction. To train the eApp, we run a simulation in which EFL begins with an untrained mobility prediction model (a deep neural network with 4 layers). The non-RT system descriptor periodically – every 10 seconds – collects the clients’ locations as they move, storing these positions in the clients’ descriptors. This historical movement data is used to train the eApp, which uses past clients’ positions to predict their future locations. The loss function of the neural network is defined as the mean squared error between the predicted and actual positions. Fig. 3 shows the convergence of the loss function and the actual vs predicted trajectory of a client. Following the arrow numbered 4 in Fig. 2, the client recruitment for each FL-service is conducted based on the dataset type of each client (16 clients with MNIST dataset are assigned to FL-service 1 and 16 clients with Fashion MNIST dataset are assigned to FL-service 2). Referring to Fig. 2, following the arrow number 5, the trained eApp is then used as a component in the near-RT FL controller to predict the system state, based on which network control managements are performed as follows.
@@ III-A2 Near-RT FL controller
Referring to Fig. 2, following the arrow number 5, we implement functionalities of near-RT FL controller for slice creation, elastic slice scaling, and mobility management jointly as an optimization problem. We define the operational time window for the near-RT FL controller as 3 seconds: every 3 seconds, near-RT FL controller solves the optimization problem to adjust service-level resources provisioning (scale down/up) for each slice.
This optimization problem aims to determine the optimal RAN slice configuration for FL-services at each O-RU. Specifically, the near-RT FL controller leverages the eApp to predict client locations over the next 3 seconds, based on which it estimates the perspective-driven distribution of clients in terms of dataset types. It then estimates the data rate between clients and O-RUs. We calculate the data rates similar to [9] as a function of the distance between clients and O-RUs. The objective function of this optimization is the minimization of $a_{1}\delta_{1}+a_{2}\delta_{2}$, which is a weighted sum of the total latency of model transmission from the clients of FL-service 1 ($\delta_{1}$) and those of FL-service 2 ($\delta_{1}$) per each O-RU. Motivated by a stricter latency requirement of FL-service 2, we form this weighted sum by coefficients $a_{1}=30$ and $a_{2}=100$. This optimization obtains the optimal client to O-RU assignment and split of bandwidth of each OR-U among the two FL-services.
Near-RT FL controller then creates a slice creation request for each FL-service, including the required bandwidth obtained above. It then sends the request to O-RAN near-RT RIC to create a RAN slice at each O-RU for each FL-service. Near-RT controller solves the above optimization every 3 seconds to perform slice scaling down/up (i.e., adjusting the bandwidth) and handover operations.
@@ III-A3 FL MAC scheduler
Referring to Fig. 2, following the arrow number 6, we design FL MAC schedulers fed with the information from the near-RT FL controller. Since the MAC scheduler operates in real-time (within a 10-millisecond interval), we develop a lightweight FL MAC scheduler for each slice. These schedulers use eApps to predict client locations over the next 10 milliseconds. Based on this prediction, they calculate the channel gains between clients and their connected O-RU. To decrease the number of unsuccessful clients of FL-services (which is a service-level performance metric), the FL MAC schedulers first sort the clients based on their channel gains, and then allocate enough bandwidth to each client (starting from the client with the highest channel gain) to satisfy the FL-services delay constraints.
III-B Optimization Setup
We compare EFL with the following baseline methods.
(Baseline 1) Conventional FL Methods [3, 13, 2]
This baseline reflects traditional FL methods and conducts snapshot-based analysis of the system with ignoring the unique characteristics of O-RAN. It solves a similar optimization problem to the one described in near-RT FL controller above but – since O-RAN functionalities are not considered – this problem obtains client-level bandwidth allocation and is solved at the beginning of each global round.
(Baseline 2) Network Functionalities for Native Wireless Services [4, 5, 14]
This baseline represents existing network control functionalities used for native wireless services by solving an optimization problem, the objective of which is allocating bandwidth to clients to maximize each client’s data rate. This baseline does not incorporate service-level objectives or constraints of FL-services.
To solve all the optimizations above, we apply techniques in our previous works [13], which involve transforming the optimization problems into convex forms and solving them using CVXPY [15]. Below, we discuss the simulation results.
III-C Simulation Results and Discussions
The left diagrams of both top and bottom plots in Fig. 4 show that Baseline 1 diminishes the model accuracy (around $40\%$) for FL-service 2. This occurs because Baseline 1 allocates resources at the start of each global round. As a result, the initially optimal resource allocation becomes suboptimal due to changes in the network caused by client movement, leading to over- or under-provisioning of resources. This, in turn, can cause delays in model transmission, preventing some clients from successfully sending their local models to O-RUs – 8 and 7 unsuccessful clients for FL-service 1 and FL-service 2, respectively. Due to the severe dataset heterogeneity of the clients in FL-service 2, this results in a greater drop in accuracy for this service.
As shown in the middle diagrams of the top and bottom plots in Fig. 4, Baseline 2 results in failure of the majority of clients. This is because, Baseline 2 focuses on maximizing each client’s data rate (a per-client performance metric), rather than maximizing the number of successful clients for FL-services (which is a service-level performance metric). This highlights how MAC-layer resource allocation and mobility management designed for native wireless services fall short in addressing the unique demands of FL-services.
As illustrated in the right plots of the top and bottom diagrams in Fig. 4, EFL outperforms both baselines. Our approach improves model accuracy of both FL-services ($90\%$ for FL-service 1 and $86\%$ for FL-service 2) as more clients can transmit their local models to the O-RUs.
To further highlight the importance of service-level performance requirements, Fig. 5 presents an ablation study showing how the weight/importance assigned to FL-service 2 (i.e., $a_{2}$ in the objective function) affects the overall performance of both FL-services. It can be seen that an increase in the importance of FL-service 2 increases its number of successful clients – since the network resources (i.e., bandwidth) will be given to FL-service 2 with a higher priority – while decreasing the number of successful clients of FL-service 1.

### v6

@@ II-F System Operation
II-G Challenges of Integration into the Existing O-RAN Implementations
Integrating EFL into existing commercial O-RAN implementations such as the OpenAirInterface (OAI) and Software-Defined RAN (SD-RAN) [8] poses several challenges due to their current design limitations. First, both OAI and SD-RAN provide limited built-in support for FL-specific service-level objectives, such as model convergence time, client heterogeneity, and multi-service orchestration. Extending them to support our EFL framework would require custom modifications to their RIC platforms to support estimator applications (eApps) and FL-specific slicing logic. Second, achieving real-time responsiveness across non-RT, near-RT, and RT layers necessitates tight synchronization and lightweight cross-layer APIs, which may not be natively supported in current RIC/xApp/rApp pipelines. For example, deploying FL MAC schedulers at the O-DU level would require modification of the real-time scheduler in OAI or SD-RAN’s platform to accommodate FL client-specific metrics such as local model delivery deadlines. Third, interoperability between vendor-specific O-RUs and generalized FL slice descriptors remains a challenge due to heterogeneous support for open fronthaul. Despite these challenges, the modular nature of OAI and SD-RAN makes them promising platforms for incremental integration of EFL.
@@ III Case Study and Prototyping
For prototyping, we consider a dense urban O-RAN scenario with vehicular clients in Porto city. The positions and mobility trajectories of the clients are drawn from the real-world Porto city taxi dataset: https://www.kaggle.com/datasets/crailtap/taxi-trajectory/data. We model an O-RAN deployment with 4 O-DUs, 18 O-RUs, and 150 FL clients (see Fig. 3). The O-RUs provide coverage over a hexagonal grid with a radius of 500 m, each assumed to have 5 MHz of available bandwidth. We also assume that FL clients have a maximum transmit power of 300 mW. The O-RAN environment, including signal forms, channels, and data rates, is modeled according to [9]. It is worth noting that the optimization problems in our framework are solved independently for each O-DU and O-RU in parallel, enabling scalability to accommodate large numbers of clients.
We consider three FL-services, each recruiting $50$ clients, which utilize convolutional neural networks (CNNs) to train on CIFAR10 (FL-service 1), Fashion-MNIST (FL-service 2), and MNIST (FL-service 3) datasets [10]. In FL-service 1, data points are distributed based on Dirichlet distribution $\mathsf{Dir}(\alpha)$ [11] with parameter $\alpha{=}0.1$, making clients’ datasets highly heterogeneous. For FL-service 2 and FL-service 3, we use $\alpha{=}10$. Each FL service runs for 100 global training rounds, where global aggregations are performed using the FedAvg method [12]. We assume that FL-service 1 has a strict delay constraint of $15$ seconds for receiving local models from its clients, while FL services 2 and 3 have a more relaxed delay constraint of $20$ seconds. Clients that fail to fully transfer their models to their O-RUs within the designated time window are considered unsuccessful.
$\star$ The details of our implementations, including all source codes and extended discussions on the mathematical formulations, are provided in our GitHub repository: https://github.com/payamsiabd/Elastic_FL_Over_O_RAN.git
Figure 5: Energy consumption and computation overhead of different methods. Our EFL method maintains a low energy consumption and computation overhead compared to the baselines with the added benefit of model performance gains as reflected in Fig. 4. The numbers that correspond to each bar are represented above it for clarity. Additional insights can be found in the annotated text within the plots.
Figure 6: Energy consumption and the number of successful clients of EFL upon varying the importance of energy efficiency $\lambda_{E}$ in the objective function of our optimization. Increasing $\lambda_{E}$ leads to energy savings (the numbers that correspond to each bar are represented above it for clarity) but it results in a drop in the number of successful clients, which in turn causes reductions in the model performance as reflected in Fig. 4. This reveals the tradeoff between energy savings and model performance. Additional insights can be found in the annotated text within the plots.
@@ III-A1 Non-RT system descriptor
To prototype this component, we first focus on arrows numbered 1 and 2 at the bottom of Fig. 2 and implement features for storing client descriptor (containing clients’ locations and dataset types) and FL-service descriptors (containing delay constraints for receiving local models). Considering the arrow numbered 3 in Fig. 2, we develop an eApp for mobility pattern prediction. To train the eApp, we run a simulation where EFL starts with an untrained mobility prediction model, which uses past clients’ positions to predict their future locations for the next 0.01 seconds (the model architecture is provided in the above GitHub). The model optimizes the mean squared error (MSE) loss function to minimize prediction errors in mobility patterns. We note that our eApps train rapidly (within a few learning iterations; see our GitHub for details) and provide accurate location predictions for the next 0.01 seconds. As a result, the impact of estimation errors is negligible, making our optimizations robust to such inaccuracies.
Following the arrow numbered 4 in Fig. 2, the client recruitment for each FL-service is conducted based on the dataset type of each client (50 clients with CIFAR10 dataset are assigned to FL-service 1, 50 clients with Fashion-MNIST dataset are assigned to FL-service 2, and 50 clients with MNIST dataset are assigned to FL-service 3). Referring to Fig. 2, following the arrow number 5, the trained eApp is then used as a component in the near-RT FL controller to predict the system state, based on which network control managements are performed as follows.
@@ III-A2 Near-RT FL controller
Referring to Fig. 2, following the arrow number 5, we implement functionalities of near-RT FL controller for slice and mobility management jointly as an optimization problem. We define the operational time window for the near-RT FL controller as 400 milliseconds: every 400 milliseconds, the near-RT FL controller first leverages the eApp to predict client locations for the next 10 milliseconds and then estimates the data rate between clients and O-RUs. Near-RT FL controller then solves an optimization problem to determine the optimal transmit power of clients and the bandwidth allocated to each client to maximize $\sum_{s=1}^{3}\delta_{s}-\lambda_{E}\sum_{u}P_{u}$, which represents the difference between two terms: (i) an FL-service–level metric, formulated as the sum of successful clients $\delta_{s}$ of each FL-service $s$, and (ii) the total energy consumption of all clients $\sum_{u}P_{u}$. Here, $\lambda_{E}$ specifies the importance assigned to energy efficiency, and $P_{u}$ denotes the transmit power of client $u$. Detailed formulation of our optimization problem are provided in the above GitHub.
Near-RT FL controller then creates a slice creation request for each FL-service that specifies (i) the maximum transmit power allowed for each client within the slice and (ii) the set of FL clients connected to each O-RU, defined as those whose allocated power exceeds a small threshold of $0.01$ W. It then sends the request to O-RAN near-RT RIC to create a RAN slice at each O-RU for each FL-service.
@@ III-A3 FL MAC scheduler
Referring to Fig. 2, and following arrow number 6, we design FL MAC schedulers that operate in real-time (every 10 milliseconds) using information from the near-RT FL controller. The MAC schedulers leverage eApps to predict client locations over the next 10 milliseconds and, based on these predictions, estimate the channel gains between clients and their connected O-RUs. They then solve lightweight optimization problems to allocate bandwidth and power to clients. These optimization problems follow the same structure as the one solved by the near-RT controller but differ in two key ways: (i) they are solved per O-RU and only over the clients currently connected to that O-RU (as determined by the near-RT controller), and (ii) the client power on each link is upper bounded by the slice-level power cap computed by the near-RT controller.
We compare EFL against five methods:
  1. 1.
    Single time-scale dynamic handover and MAC scheduling (ST-HMAC): both handover and MAC scheduling are performed at each MAC scheduling interval (i.e., every 0.01 s), inspired by [13].
  2. 2.
    Snapshot handover and dynamic MAC scheduling (S-HMAC): handover is performed only once at the start of each global round, while MAC scheduling is updated dynamically every 0.01 s, inspired by [14].
  3. 3.
    Snapshot handover and MAC scheduling (SH-MAC): inspired by conventional FL orchestrators [3, 12, 2], this method performs both handover and MAC scheduling once at the beginning of each global round.
  4. 4.
    General MAC scheduling (G-MAC): resembling current MAC schedulers used in general internet services such as video streaming [4, 5, 15], this method does not incorporate FL-specific QoS considerations such as the number of successful clients or prioritization among FL services.
  5. 5.
    Full participation upper bound (FP-UP): an idealized benchmark where all FL clients in all services are always able to successfully transmit their local models to the server, serving as a hypothetical upper bound on test accuracy (due to violating the resource constraints).
III-B Results and Discussions
The top plot in Fig. 4 shows that ST-HMAC yields the worst performance as it performs handover operations and MAC scheduling every 0.01 seconds, which introduces substantial overhead. Specifically, since FL clients must upload their local models to the server within a strict deadline, the time wasted by these frequent control operations reduces the available transmission window, thereby increasing the number of unsuccessful clients. This is further reflected by its worst performance across the methods in the bottom plot of Fig. 4. Also, it can be seen that S-HMAC performs better than the other baselines but worse than our EFL. This is because, although S-HMAC carries out dynamic MAC scheduling, it still performs handover operations only at the start of each global round (the initially optimal connections between clients and O-RUs may become suboptimal due to client mobility). Moreover, G-MAC results in more failure of clients compared to EFL as it focuses on maximizing each client’s data rate (i.e., a per-client performance metric), rather than maximizing the number of successful clients for FL-services (i.e., service-level performance metric). This highlights how MAC-layer resource allocation and mobility management designed for native wireless services fall short in addressing the unique demands of FL-services. As illustrated in Fig. 4, our EFL outperforms all baselines by enabling more clients to successfully transmit their local models to O-RUs. This is further reflected by its resulting model accuracies that are comparable to the upper bound achieved by FP-UP.
The left plot of Fig. 5 depicts the energy consumption of all methods. G-MAC consumes the most energy as its objective prioritizes maximizing user data rates, which in turn drives higher power consumption. The energy consumption of EFL is comparable to ST-HMAC, S-HMAC, and SH-MAC, demonstrating that our framework does not impose additional energy burdens. Importantly, while maintaining similar energy costs, EFL yields the highest model accuracy among all algorithms, as discussed above, by enabling the largest number of successful clients to transmit their models within the deadline. Also, the right plot in Fig. 5 reports the average computation time overhead per global round incurred by handover operations and MAC scheduling. EFL achieves a computation overhead that is close to the best baselines; however, it is worth noting that S-HMAC and SH-HMAC yield lower computation overheads compared to EFL as they perform handover operation only once at the beginning of each global round. Nevertheless, these additional overheads of EFL remain tractable and are justified by the gains in the previously discussed client success rate and model accuracy. This highlights the advantage of our design in balancing computational overhead, energy efficiency, and FL accuracy.
In Fig. 6, we conduct analyses on the value of $\lambda_{E}$, which tunes the importance of energy term in our optimization, to show how varying this coefficient affects both the success rate of clients and the total energy consumption of EFL. As shown, by increasing the value of $\lambda_{E}$, our method prioritizes energy savings (i.e., it reduces the transmit power of clients) by sacrificing the number successful clients, which will lead to a drop in model accuracy as we studied in Fig. 4. These results highlight the tradeoff between energy conservation and client participation, offering insights into how EFL can be tailored to energy-constrained IoT devices depending on the prioritized objective (model performance vs. energy efficiency).